OpenStudy (anonymous):

simplify using only positive exponents: (4x^3 y^-2)^0/(2x^-4 y^3)^-3 (3x^5 y)^-2

OpenStudy (anonymous):

can you draw it?

zepdrix (zepdrix):

\[\huge \frac{(4x^3y^{-2})^0}{(2x^-4y^3)^{-3}(3x^5y)^{-2}}\] It's a little hard to read, is this what the problem looks like?

OpenStudy (anonymous):

yes except its 2x^-4

OpenStudy (anonymous):

the numerator is just 1 :P

OpenStudy (anonymous):

because its raised to the 0 power right?

zepdrix (zepdrix):

\[\huge \frac{(4x^3y^{-2})^0}{(2x^{-4}y^3)^{-3}(3x^5y)^{-2}}\]Ah sorry bout that :) Formatting error.

OpenStudy (anonymous):

no worries! :]

OpenStudy (anonymous):

thats awesome how you did that, and clear text!

OpenStudy (anonymous):

yup laylam

zepdrix (zepdrix):

Yes because of the 0 power. When you square something, you multiply it by itself, so it becomes squared. When you raise something to the 1st power, you do nothing to it. When you raise something to the 0th power, you divide it by itself, giving you 1.

zepdrix (zepdrix):

\[\huge \frac{1}{(2x^{-4}y^3)^{-3}(3x^5y)^{-2}}\]Just need to take care of the mess on the bottom now :)

zepdrix (zepdrix):

The negative power tells us that we can FLIP the term if we change it to positive. Example:\[\huge \frac{1}{x^{-2}} \qquad = \qquad x^2\]See how I changed it to a positive and brought it up top?

OpenStudy (anonymous):

okay great! So since (2x^-4 y^3)^-3 is raised to the -3, do we cube that?

OpenStudy (anonymous):

yup!

zepdrix (zepdrix):

We want to first bring it up top, since it's a negative in the denominator :) But then, yes, we'll be able to cube it after that.

OpenStudy (anonymous):

okay so it would be (2x^-4 y^3)^3 (3x^5y)^2 at the top?

zepdrix (zepdrix):

Yes good good c:

OpenStudy (anonymous):

yay!

zepdrix (zepdrix):

Remember how to apply the exponents to terms in a bracket set? You will apply it to EVERYTHING in the brackets!

OpenStudy (anonymous):

so it would be (2x^-12 y^9)^3 (3x^10y)^2

zepdrix (zepdrix):

Well, a couple things to be careful about. When you apply the exponent, you shouldn't have it floating outside anymore. Because the way you have it written now, it looks like you still need to apply the exponent to it. Also, don't forget to apply the exponent to CONSTANT terms, not just variables. (Hint hint: the 2 and 3)

OpenStudy (anonymous):

oh I see, so like this? (6x^-12 3y^9) (6x^10 2y)

zepdrix (zepdrix):

From this,\[\huge (2x^{-4}y^3)^{3}(3x^5y)^{2}\] You got, \[\huge (\color{orangered}{6}x^{-12} \color{orangered}{3}y^9)(\color{orangered}{6}x^{10}\color{orangered}{y})\] The red are the parts we need to fix.

zepdrix (zepdrix):

The constant in the first bracket should be,\[\huge 2^3\]Which is not 6 :O

OpenStudy (anonymous):

is the x still there?

zepdrix (zepdrix):

Anything I left in black was correct, so don't worry about the X's :)

OpenStudy (anonymous):

are you just attaching the exponent to each variable and not multiplying it?

zepdrix (zepdrix):

\[\huge (2x^{-4}y^3)^3 \qquad = \qquad 2^3(x^{-4})^3(y^3)^3\]We cube eeeeeverything. See how that works? :D

OpenStudy (anonymous):

I'm confused :/ then for (y^3)^3 why isn't it going to be (3y^9)?

zepdrix (zepdrix):

If you wanted to, you could think of the y term like this,\[\huge (1\cdot y^3)^3\]When we cube this, we'll get,\[\huge 1^3\cdot y^9 \quad =\quad y^9\] I'm not really sure where your 3 is coming from.

OpenStudy (anonymous):

ohh I see! I just multiplied the exponent to the y

OpenStudy (anonymous):

so would it be (2x^-12 y^9) (3x^10 y^2)

zepdrix (zepdrix):

You have your x's and y's correct. The constants we still need to fix though.

zepdrix (zepdrix):

You didn't apply the exponent to your constants in front.

OpenStudy (anonymous):

(2^3 x^-12 y^9) (3^2 x^10 y^2)

zepdrix (zepdrix):

Yes good :) now simplify the constants down silly! heh

OpenStudy (anonymous):

ah see thats where I was confused when you said "apply the exponent to the constant" haha thats where I multiplied -_-

zepdrix (zepdrix):

Hah XD my bad, maybe i could have explained that better.

OpenStudy (anonymous):

no worries! Okay so it's (8 x^-12 y^9)(9 x^10 y^2)

zepdrix (zepdrix):

Yes looks good. and from here we can simplify further I'm afraid!! :(

OpenStudy (anonymous):

:o

zepdrix (zepdrix):

We can multiply out the brackets :OOO

OpenStudy (anonymous):

multiplying the x's and the y's together?

zepdrix (zepdrix):

Luckily, each bracket has only one TERM, so no FOIL nonsense.

zepdrix (zepdrix):

Yes :) Remember what to do with the exponents when you multiply bases?

OpenStudy (anonymous):

phew!

OpenStudy (anonymous):

they get added!

zepdrix (zepdrix):

Yesss c:

OpenStudy (anonymous):

72x^-3 2y^11 !!!!

zepdrix (zepdrix):

Uh ohhhhhhhhhhhhhhhhhhhhhhh,\[\huge x^{-12}\cdot x^{10}=x^{\color{orangered}{-3}}\]

zepdrix (zepdrix):

Red means bad! XD lol

OpenStudy (anonymous):

but what about 8*9?!

zepdrix (zepdrix):

\[\huge 8 \cdot 9=72\]Ok that part looks good. Why is there a 2 in front of the y? :\

OpenStudy (anonymous):

so thats right 72x^-3? oh I thought because there are two y's :/ so then its just

OpenStudy (anonymous):

72x^-3 y^11 ?

zepdrix (zepdrix):

Why are you getting -3 on the X?

OpenStudy (anonymous):

-2!! my bad lol

zepdrix (zepdrix):

Yay team \:D/

OpenStudy (anonymous):

omg haha silly mistakes!

OpenStudy (anonymous):

but thank you so much!!!

OpenStudy (anonymous):

:D

zepdrix (zepdrix):

Yah easy to make silly mistakes :D long problem. heh